OpenStudy (anonymous):

Which of the following points is a solution to the graph of the line, y = 4x - 3? A. (-11, -2) B. (5, 2) C. (-3, 0) D. (3, 9)

OpenStudy (anonymous):

D.(3,9)

OpenStudy (e.cociuba):

Find x: y=4x-3 x=1 y=4(1)-3 y=1 ----------- x=0 Plug 0 back into the equation to get y: y=4(0)-3 y=-3 =(-3,0)